# Paraphrase Types Elicit Prompt Engineering Capabilities

### v1

@@ Abstract
Much of the success of modern language models depends on finding a suitable prompt to instruct the model. Until now, it has been largely unknown how variations in the linguistic expression of prompts affect these models. This study systematically and empirically evaluates which linguistic features influence models through paraphrase types, i.e., different linguistic changes at particular positions. We measure behavioral changes for five models across 120 tasks and six families of paraphrases (i.e., morphology, syntax, lexicon, lexico-syntax, discourse, and others). We also control for other prompt engineering factors (e.g., prompt length, lexical diversity, and proximity to training data). Our results show a potential for language models to improve tasks when their prompts are adapted in specific paraphrase types (e.g., 6.7% median gain in Mixtral 8x7B; 5.5% in LLaMA 3 8B; cf. Figure 1). In particular, changes in morphology and lexicon, i.e., the vocabulary used, showed promise in improving prompts. These findings contribute to developing more robust language models capable of handling variability in linguistic expression. Code: https://github.com/jpwahle/prompt-paraphrase.
@@ 2 Related Work
Prompt-based learning has become a new trend, where pre-trained language models perform prediction using a template description of the intended task, and the model derives the necessary information without the need for gradient updates Liu et al. (2023a). Prompt-based learning has many advantages over traditional supervised learning, as the same pre-trained LLM can perform different unseen tasks without fine-tuning. Early prompt tuning methods focused on integrating trainable continuous prompt embeddings to perform various NLU tasks Gu et al. (2022); Liu et al. (2021, 2023b). Discrete prompts representing actual tokens have become more popular than continuous embeddings in prompt engineering, arguably because they are more accessible and linguistically interpretable.
Discrete methods used cloze-style phrases and differentiable prompts to improve few-shot learning capabilities Schick and Schütze (2021); Zhang et al. (2022). Shin et al. (2020) proposed AutoPrompt, which uses a masked language model to find variations of prompts and has been used for various downstream tasks such as plagiarism detection Wahle et al. (2022). Zhou et al. (2022) introduced Automatic Prompt Engineer, which searches over a pool of prompt candidates proposed by a second LLM. Other work focused on generating prompts for knowledge extraction from LLMs Jiang et al. (2020) or investigated instruction induction and model tuning with minimal human intervention Honovich et al. (2022b, a). Recent work uses variations of Self-Instruct to improve model prompting Wang et al. (2022a), but the prompt aspects leading to better results are largely unknown.
Reynolds and McDonell (2021); Hu and Levy (2023) evaluate the role of prompts in model control and the limitations of metalinguistic prompts for assessing linguistic knowledge in LLMs. Studies by Leidinger et al. (2023) and Lu et al. (2023) examined the influence of linguistic features of prompts on LLM performance with hand-selected prompt variations (e.g., tense, modality). Sorensen et al. (2022); Yang et al. (2023) explored the optimization of prompt selection using unified scoring frameworks and unsupervised techniques based on mutual information.
@@ 4 Experiments
Ans. We start with a seed prompt designed for a specific task, e.g., “In this task, you are given Twitter posts. Your task is to label the post’s emotion as sadness, joy, disgust, anger, fear, or surprise.” We paraphrase each prompt with one of 26 linguistic changes (e.g., syntax, lexicon) using the fine-tuned model trained by Wahle et al. (2023). See Table 7 and Figures 8 and 10 in Appendix A for an extensive list of types, full prompt template and examples respectively. For the original prompt and its 26 paraphrases, we test each model’s few-shot settings. Depending on the task, the number of few-shot examples ranges from 3 to 6 (sum of positive and negative examples).
@@
Specific perturbations of the prompt affect the performance of models on tasks from varying domains differently. Morphology changes in the prompt have the largest gains in wrong candidate generation (+26.0%), question generation (+21.5%), and textual entailment (+17.5%). Lexical changes show consistent gains in summarization (10.8%), wrong candidate generation (+8.9%), and title generation (+7.1%). These tasks rely on semantic precision and vocabulary richness to interpret or generate nuanced responses. Positive effects with discourse changes can be observed in tasks that require an understanding of longer contexts or multi-sentence structures, such as summarization (+7.8%). Detailed results for each model are available in Tables 8, 10, 11, 9, 13 and 12 in Appendix C.
@@ 5 Conclusion
Our results show that the same paraphrase changes can potentially improve results for one model but harm another. The same is true for different changes in the same task. In a task like sentiment, sometimes the same polarity substitutions lead to improvements, and sometimes they do not. Variance across examples and models does play a role here but another possible reason is that models trained to generate paraphrase types only have a certain accuracy, leading to models sometimes confusing one type for another or generating an incorrect type. However, at a large aggregate level (3.24 million prompt-example combinations), our results provide important trends about prompts despite these margins of error. Better models for the controlled paraphrase generation will provide more accurate results in the future.
@@ Appendix A Details on the Prompts & Tasks
Prompts. We use the template shown in Figure 8 to construct prompts for each of the 120 tasks. In Figure 10, we show an example for the numerical commonsense reasoning task NumerSense in which the model has to predict a blank token ‘_’ Lin et al. (2020). To paraphrase prompts, we use one of 26 paraphrase types in six groups shown in Table 7.
Tasks. We provide an overview of all 120 tasks together with the paraphrase type that, when applied to prompts in that task domain, had the most positive influence together with the potential downstream task performance gain over the original prompt in Tables 2, 3, 4 and 5. Further, we want to note an observation of our experiments showing that tasks containing content judged as unethical (such as toxic language detection on Twitter) led to many denied responses by the chat or instruction-finetuned models. Likely because of their post-training, models answered that they could not respond to these questions, although we did not ask them to produce new harmful content but to classify existing content. We removed these tasks from our final analysis. This observation raises questions about how many powerful models can be used in the future to classify toxic content when their guardrails prevent them from answering.
@@ Appendix C Additional Results
Prompt Complexity Correlations. Table 1 shows the individual correlations between the three markers of prompt complexity and downstream task performance. All correlations are close to zero, only very small correlations exist for question answering (0.13 - 0.14) and sentence composition (0.10 - 0.11). We could generally find no evidence that the complexity of the prompt contributes to gains or losses in performance outcomes.
Results by Task and Type. We show a detailed table of avgerage downstream task performance gain or loss over the original prompt of a task when paraphrasing the prompt using modifications of one of the six different paraphrase groups in Table 8. We decompose this table for each model in Tables 10, 9, 11, 13 and 12.
@@
Task Family Morphology Syntax Lexicon Lex.-Syn. Discourse Others
Answerability Classification +1.7% 0.0% 0.0% 0.0% 0.0% 0.0%
Commonsense Classification +3.6% $\uparrow$ 0.0% 0.0% 0.0% 0.0% 0.0%
Coreference Resolution +1.1% +2.2% 0.0% 0.0% 0.0% 0.0%
Dialogue Generation +4.7% $\uparrow$ +2.2% +4.1% $\uparrow$ +0.5% +0.6% +3.0% $\uparrow$
Fill in The Blank +4.8% $\uparrow$ 0.0% 0.0% 0.0% 0.0% 0.0%
Information Extraction +2.2% -8.7% $\downarrow$ -3.3% $\downarrow$ -0.8% +0.8% +3.4% $\uparrow$
Named Entity Recognition +3.2% $\uparrow$ 0.0% -0.5% 0.0% 0.0% 0.0%
Program Execution -6.0% $\downarrow$ +1.5% -4.6% $\downarrow$ +5.1% $\uparrow$ 0.0% 0.0%
Question Answering +4.3% $\uparrow$ +5.6% $\uparrow$ -7.1% $\downarrow$ +10.3% $\uparrow$ +9.9% $\uparrow$ +30.4% $\uparrow$
Question Generation +12.3% $\uparrow$ +7.8% $\uparrow$ -8.1% $\downarrow$ +6.5% $\uparrow$ +10.2% $\uparrow$ -3.0%
Question Rewriting -4.4% $\downarrow$ -53.7% $\downarrow$ -5.7% $\downarrow$ -13.4% $\downarrow$ -4.8% $\downarrow$ -4.6% $\downarrow$
Question Understanding +9.3% $\uparrow$ 0.0% 0.0% 0.0% 0.0% 0.0%
Sentence Composition -8.6% $\downarrow$ +2.6% +2.8% -0.9% -8.7% $\downarrow$ +10.6% $\uparrow$
Sentiment Analysis +2.2% 0.0% 0.0% 0.0% 0.0% 0.0%
Summarization +11.7% $\uparrow$ +9.9% $\uparrow$ +8.1% $\uparrow$ +7.7% $\uparrow$ -1.8% +3.6% $\uparrow$
Text Categorization +1.5% -6.3% $\downarrow$ -0.8% +20.0% $\uparrow$ 0.0% 0.0%
Text Completion +2.7% -1.1% -4.4% $\downarrow$ -0.2% +15.2% $\uparrow$ -0.2%
Text Matching +1.0% 0.0% 0.0% 0.0% 0.0% 0.0%
Text Quality Evaluation +4.1% $\uparrow$ 0.0% 0.0% 0.0% 0.0% 0.0%
Text to Code +1.9% -4.9% $\downarrow$ -2.0% -2.4% 0.0% -1.4%
Textual Entailment +3.8% $\uparrow$ 0.0% 0.0% 0.0% 0.0% 0.0%
Title Generation +7.4% $\uparrow$ +4.1% $\uparrow$ +7.3% $\uparrow$ -7.5% $\downarrow$ +5.0% $\uparrow$ 0.0%
Word Semantics +2.5% 0.0% 0.0% 0.0% 0.0% 0.0%
Wrong Candidate Generation +26.8% $\uparrow$ -6.6% $\downarrow$ +8.0% $\uparrow$ -4.7% $\downarrow$ +1.6% +21.4% $\uparrow$

### v4

# EMNLP 2024 Paraphrase Types Elicit Prompt Engineering Capabilities

@@ Abstract
Much of the success of modern language models depends on finding a suitable prompt to instruct the model. Until now, it has been largely unknown how variations in the linguistic expression of prompts affect these models. This study systematically and empirically evaluates which linguistic features influence models through paraphrase types, i.e., different linguistic changes at particular positions. We measure behavioral changes for five models across 120 tasks and six families of paraphrases (i.e., morphology, syntax, lexicon, lexico-syntax, discourse, and others). We also control for other prompt engineering factors (e.g., prompt length, lexical diversity, and proximity to training data). Our results show a potential for language models to improve tasks when their prompts are adapted in specific paraphrase types (e.g., 6.7% median gain in Mixtral 8x7B; 5.5% in LLaMA 3 8B; cf. Figure 1). In particular, changes in morphology and lexicon, i.e., the vocabulary used, showed promise in improving prompts. These findings contribute to developing more robust language models capable of handling variability in linguistic expression. Code: https://github.com/jpwahle/emnlp24-prompt-paraphrase.
@@ 2 Related Work
Prompt-based learning has become a new trend, where pre-trained language models perform prediction using a template description of the intended task, and the model derives the necessary information without the need for gradient updates Liu et al. (2023a). Prompt-based learning has advantages over traditional supervised learning, as the same pre-trained LLM can perform different unseen tasks without fine-tuning. Early prompt tuning methods focused on integrating trainable continuous prompt embeddings to perform various NLU tasks Gu et al. (2022); Liu et al. (2021, 2023b). Discrete prompts representing actual tokens have become more popular than continuous embeddings in prompt engineering, arguably because they are more accessible and linguistically interpretable.
Discrete methods used cloze-style phrases and differentiable prompts to improve few-shot learning capabilities Schick and Schütze (2021); Zhang et al. (2022). Shin et al. (2020) proposed AutoPrompt, which uses a masked language model to find variations of prompts and has been used for various downstream tasks such as plagiarism detection Wahle et al. (2022). Zhou et al. (2022) introduced Automatic Prompt Engineer, which searches over a pool of prompt candidates proposed by a second LLM. Other work focused on generating prompts for knowledge extraction from LLMs Jiang et al. (2020) or investigated instruction induction and model tuning with minimal human intervention Honovich et al. (2022b, a). Recent work uses variations of Self-Instruct to improve model prompting Wang et al. (2022a).
Reynolds and McDonell (2021); Hu and Levy (2023) evaluate the role of prompts in model control and the limitations of metalinguistic prompts for assessing linguistic knowledge in LLMs. Studies by Leidinger et al. (2023) and Lu et al. (2023) examined the influence of linguistic features of prompts on LLM performance with hand-selected prompt variations (e.g., tense, modality). Mizrahi et al. (2024) demonstrated through empirical analysis that LLM performance can vary based on how task instructions are phrased. Sorensen et al. (2022); Yang et al. (2023) explored the optimization of prompt selection using unified scoring frameworks and unsupervised techniques based on mutual information.
@@ 4 Experiments
Ans. We start with a seed prompt designed for a specific task, e.g., “In this task, you are given Twitter posts. Your task is to label the post’s emotion as sadness, joy, disgust, anger, fear, or surprise.” We paraphrase each prompt with one of 26 linguistic changes (e.g., syntax, lexicon) using the gpt-3.5-turbo-16k model, which we fine-tuned in the same way as Wahle et al. (2023) using the ETPC dataset Kovatchev et al. (2018). See Table 9 and Figures 8 and 10 in Appendix A for an extensive list of types, full prompt template and examples respectively. For the original prompt and its 26 paraphrases, we test each model’s few-shot settings. Depending on the task, the number of few-shot examples ranges from 3 to 6 (sum of positive and negative examples).
@@
Specific perturbations of the prompt affect the performance of models on tasks from varying domains differently. Morphology changes in the prompt have the largest gains in wrong candidate generation (+26.0%), question generation (+21.5%), and textual entailment (+17.5%). Lexical changes show consistent gains in summarization (10.8%), wrong candidate generation (+8.9%), and title generation (+7.1%). These tasks rely on semantic precision and vocabulary richness to interpret or generate nuanced responses. Positive effects with discourse changes can be observed in tasks that require an understanding of longer contexts or multi-sentence structures, such as summarization (+7.8%). Detailed results for each model are available in Tables 10, 12, 13, 11, 15 and 14 in Appendix C.
@@ 5 Conclusion
Our results show that the same paraphrase changes can potentially improve results for one model but harm another. The same is true for different changes in the same task. In a task like sentiment, sometimes the same polarity substitutions lead to improvements, and sometimes they do not. Variance across examples and models does play a role here but another possible reason is that models trained to generate paraphrase types only have a certain accuracy, leading to models sometimes confusing one type for another or generating an incorrect type as recent work shows Meier et al. (2024). However, at a large aggregate level (3.24 million prompt-example combinations), our results provide important trends about prompts despite these margins of error. Better models for the controlled paraphrase generation will provide more accurate results in the future.
@@
Acknowledgements
This work was partially supported by the Lower Saxony Ministry of Science and Culture and the VW Foundation. Many thanks to Andreas Stephan and Lars Käsberg for their thoughtful discussions.
@@ Appendix A Details on the Prompts & Tasks
Prompts. We use the template shown in Figure 8 to construct prompts for each of the 120 tasks. In Figure 10, we show an example for the numerical commonsense reasoning task NumerSense in which the model has to predict a blank token ‘_’ Lin et al. (2020). To paraphrase prompts, we use one of 26 paraphrase types in six groups shown in Table 9. Five examples of paraphrases applied to prompts can be seen in Table 1. In the first case, the added specification made the instruction more precise; in the second case, the model removed seemingly unnecessary context. The third case involved replacing words or phrases with synonyms or near-synonyms that have the same meaning in the given context, while the fourth case involved restructuring the sentence or changing the way the information is presented while maintaining the same overall meaning. Finally, the fifth case replaced a single word with a phrase (or vice versa) that conveys the same meaning. Further paraphrase type definitions and perturbation examples can be found in Vila et al. (2014); Meier et al. (2024).
Tasks. We provide an overview of all 120 tasks together with the paraphrase type that, when applied to prompts in that task domain, had the most positive influence together with the potential downstream task performance gain over the original prompt in Tables 4, 5, 6 and 7. Further, we want to note an observation of our experiments showing that tasks containing content judged as unethical (such as toxic language detection on Twitter) led to many denied responses by the chat or instruction-finetuned models. Likely because of their post-training, models answered that they could not respond to these questions, although we did not ask them to produce new harmful content but to classify existing content. We removed these tasks from our final analysis. This observation raises questions about how many powerful models can be used in the future to classify toxic content when their guardrails prevent them from answering.
@@
Instruction Type Instruction Text ROUGE-L
Example 1: Addition/Deletion
Original Your task is to generate a headline (title) for this article. 0.60
Paraphrased Your task is to generate a concise headline (title) for this article. 0.78
Example 2: Ellipsis
Original Your task is to generate a headline (title) for this article. 0.60
Paraphrased Generate a headline (title) for this article. 0.67
Example 3: Same Polarity Substitution (contextual)
Original Summarize the main points of the given text in 3-4 sentences. 0.55
Paraphrased Condense the key ideas of the provided passage into 3-4 sentences. 0.72
Example 4: Syntax/discourse structure changes
Original Explain the concept of gravity as if you were teaching a 10-year-old child. 0.63
Paraphrased Imagine you’re teaching a 10-year-old child. Explain the concept of gravity. 0.71
Example 5: Synthetic/analytic substitution
Original Translate the following English text into French. 0.58
Paraphrased Convert the given English passage to its French equivalent. 0.70
Table 1: Comparison of original and paraphrased instructions for five types with downstream ROUGE-L scores.
@@ B.5 Potential Downstream Gain
Task Family Min Max Avg ($\downarrow$)
Title Generation 24 53 36.0
Fill in The Blank 25 60 43.2
Summarization 21 123 52.0
Text Quality Eval. 42 67 52.6
Named Entity Rec. 38 77 52.8
Word Semantics 39 74 53.8
Sentiment Analysis 24 115 56.6
Sentence Composition 19 84 61.8
Answerability Class. 47 70 63.6
Text Completion 40 111 66.0
Dialogue Generation 27 186 67.8
Question Answering 45 106 72.0
Program Execution 43 110 75.4
Coreference Resolution 24 219 78.8
Information Extr. 70 91 79.0
Textual Entailment 61 111 81.2
Text Categorization 55 116 88.2
Text Matching 34 203 91.0
Question Generation 29 160 102.2
Wrong Candidate Gen. 105 192 129.4
Commonsense Class. 49 202 148.8
Question Rewriting 46 753 216.0
Question Understanding 125 581 291.8
Text to Code 121 638 390.6
Overall 19 753 102.1
Table 3: The number of tokens per input prompt (computed using the GPT-2 tokenizer). Min shows the shortest prompt, Max the longest prompt, and Avg the average prompt length for all five tasks.
@@ Appendix C Additional Results
Prompt Complexity Correlations. Table 2 shows the individual correlations between the three markers of prompt complexity and downstream task performance as described in Section B.2. All correlations are close to zero, and only very small correlations exist for question answering (0.13 - 0.14) and sentence composition (0.10 - 0.11) with $p<0.05$. We could generally find no evidence that the complexity of the prompt contributes to gains or losses in performance outcomes.
Prompt Lengths. We conducted a post-hoc analysis of the number of input tokens across various task families using the GPT-2 tokenizer. As shown in Table 3, the average number of tokens varies significantly depending on the nature of the task, ranging from 36 tokens for title generation tasks to 390 tokens for text-to-code tasks. The overall distribution of input tokens is even broader when looking at individual tasks, with the shortest task, sentence composition, using as few as 19 tokens, and the longest, question rewriting, reaching up to 753 tokens. This wide spectrum of prompt lengths indicates that our study covers a diverse array of real-world use cases, from concise to extended prompts. By analyzing 120 tasks across 24 task families, we believe our conclusions hold across a broad range of prompt lengths, including those much shorter (e.g., fewer than 36 tokens) and much longer (e.g., more than 800 tokens).
Results by Task and Type. We show a detailed table of avgerage downstream task performance gain or loss over the original prompt of a task when paraphrasing the prompt using modifications of one of the six different paraphrase groups in Table 10. We decompose this table for each model in Tables 12, 11, 13, 15 and 14.
@@
Task Family Morphology Syntax Lexicon Lex.-Syn. Discourse Others Answerability Classification +1.7% 0.0% 0.0% 0.0% 0.0% 0.0% Commonsense Classification +3.6% $\uparrow$ 0.0% 0.0% 0.0% 0.0% 0.0% Coreference Resolution +1.1% +2.2% 0.0% 0.0% 0.0% 0.0% Dialogue Generation +4.7% $\uparrow$ +2.2% +4.1% $\uparrow$ +0.5% +0.6% +3.0% $\uparrow$ Fill in The Blank +4.8% $\uparrow$ 0.0% 0.0% 0.0% 0.0% 0.0% Information Extraction +2.2% -8.7% $\downarrow$ -3.3% $\downarrow$ -0.8% +0.8% +3.4% $\uparrow$ Named Entity Recognition +3.2% $\uparrow$ 0.0% -0.5% 0.0% 0.0% 0.0% Program Execution -6.0% $\downarrow$ +1.5% -4.6% $\downarrow$ +5.1% $\uparrow$ 0.0% 0.0% Question Answering +4.3% $\uparrow$ +5.6% $\uparrow$ -7.1% $\downarrow$ +10.3% $\uparrow$ +9.9% $\uparrow$ +30.4% $\uparrow$ Question Generation +12.3% $\uparrow$ +7.8% $\uparrow$ -8.1% $\downarrow$ +6.5% $\uparrow$ +10.2% $\uparrow$ -3.0% Question Rewriting -4.4% $\downarrow$ -53.7% $\downarrow$ -5.7% $\downarrow$ -13.4% $\downarrow$ -4.8% $\downarrow$ -4.6% $\downarrow$ Question Understanding +9.3% $\uparrow$ 0.0% 0.0% 0.0% 0.0% 0.0% Sentence Composition -8.6% $\downarrow$ +2.6% +2.8% -0.9% -8.7% $\downarrow$ +10.6% $\uparrow$ Sentiment Analysis +2.2% 0.0% 0.0% 0.0% 0.0% 0.0% Summarization +11.7% $\uparrow$ +9.9% $\uparrow$ +8.1% $\uparrow$ +7.7% $\uparrow$ -1.8% +3.6% $\uparrow$ Text Categorization +1.5% -6.3% $\downarrow$ -0.8% +20.0% $\uparrow$ 0.0% 0.0% Text Completion +2.7% -1.1% -4.4% $\downarrow$ -0.2% +15.2% $\uparrow$ -0.2% Text Matching +1.0% 0.0% 0.0% 0.0% 0.0% 0.0% Text Quality Evaluation +4.1% $\uparrow$ 0.0% 0.0% 0.0% 0.0% 0.0% Text to Code +1.9% -4.9% $\downarrow$ -2.0% -2.4% 0.0% -1.4% Textual Entailment +3.8% $\uparrow$ 0.0% 0.0% 0.0% 0.0% 0.0% Title Generation +7.4% $\uparrow$ +4.1% $\uparrow$ +7.3% $\uparrow$ -7.5% $\downarrow$ +5.0% $\uparrow$ 0.0% Word Semantics +2.5% 0.0% 0.0% 0.0% 0.0% 0.0% Wrong Candidate Generation +26.8% $\uparrow$ -6.6% $\downarrow$ +8.0% $\uparrow$ -4.7% $\downarrow$ +1.6% +21.4% $\uparrow$